# Systematics mitigation for catalogue-based angular power spectra

### v1

@@ Abstract
Recent work has developed a formalism for computing angular power spectra directly from catalogues containing field values at discrete positions on the sky, thereby circumventing the need to create pixelised maps of the fields, as well as avoiding aliasing and finite-resolution effects. Here, we adapt this formalism to incorporate template deprojection as a means of mitigating systematic biases in the measured angular power spectra. We also introduce and validate an alternative method of mitigating the so-called ‘deprojection bias’, caused by the loss of modes after deprojection, employing the use of simple simulations to compute a transfer function. We find that this approach performs at least as well as existing methods, and is relatively insensitive to how well one can guess the true power spectrum of the observed field, except at the largest scales ($\ell\lesssim 3$). Additionally, we develop exact expressions for the bias introduced by deprojection in the shot-noise component, which further improves the accuracy of this approach. We test our formalism on simulated datasets, demonstrating its applicability both to discretely sampled fields, and to the special case of galaxy clustering, with the survey selection function defined in terms of a random catalogue or as a continuous sky map. After computing and removing the bias in the shot noise and using a transfer function to correct for the remaining mode loss, our formalism is able to produce unbiased measurements of the angular power spectrum in all scenarios tested here. Finally, we apply our formalism to real data and demonstrate that it produces results consistent with the standard map-based pseudo-$C_{\ell}$ formalism. We implement our method in the publicly available code NaMaster.
@@ 1 Introduction
The characterisation and mitigation of systematic effects is becoming a problem of increasing relevance in cosmological surveys: as the hoards of data being collected continually increase in size, the statistical uncertainties have the potential to become comparable to those arising from systematics. A plethora of techniques have been developed to tackle this issue by attempting to quantify the dependence of the desired field on maps of the survey properties or other contaminants thought to be contributing to the observed signal. These range from linear regression techniques such as mode deprojection (also called template deprojection; Rybicki & Press, 1992; Slosar et al., 2004; Elsner et al., 2016, 2017), template subtraction (Ross et al., 2011; Ho et al., 2012), and the iterative systematics decontamination (ISD) technique employed by the Dark Energy Survey (DES; Elvin-Poole et al., 2018; Rodríguez-Monroy et al., 2022), to more sophisticated methods employing machine learning tools such as artificial neural networks (Rezaie et al., 2020) or self-organising maps (Johnston et al., 2021).

### v2

@@ Abstract
Recent work has developed a formalism for computing angular power spectra directly from catalogues containing field values at discrete positions on the sky, thereby circumventing the need to create pixelised maps of the fields, as well as avoiding aliasing and finite-resolution effects. We adapt this formalism to incorporate template deprojection for mitigating systematic biases in the measured angular power spectra. We also introduce an alternative method of mitigating the ‘deprojection bias’ – the loss of modes induced by deprojection – employing simple simulations to compute a transfer function. We find that this approach performs at least as well as existing methods, and is relatively insensitive to how well one can guess the true power spectrum of the observed field, except at the largest scales ($\ell\lesssim 3$). Additionally, we develop exact expressions for the bias introduced by deprojection in the shot-noise component, which further improves the accuracy of this approach. We test our formalism on simulated datasets, demonstrating its applicability both to discretely sampled fields, and to the special case of galaxy clustering, with the survey selection function defined in terms of a random catalogue or as a continuous sky map. After removing the bias in the shot noise and correcting for the remaining mode loss using a transfer function, our formalism produces unbiased measurements of the angular power spectrum in all scenarios tested here. Finally, we apply our formalism to real data and show it produces results consistent with the standard map-based pseudo-$C_{\ell}$ formalism. We implement our method in the public code NaMaster.
@@ 1 Introduction
The characterisation and mitigation of systematic effects is becoming a problem of increasing relevance in cosmological surveys: as the hoards of data being collected continually increase in size, the statistical uncertainties have the potential to become comparable to those arising from systematics (e.g. Ho et al., 2012a; Pullen & Hirata, 2013; Huterer et al., 2013; Agarwal et al., 2014; Shafer & Huterer, 2015; Ho et al., 2015; Weaverdyck et al., 2018). A plethora of techniques have been developed to tackle this issue by attempting to quantify the dependence of the desired field on maps of the survey properties or other contaminants thought to be contributing to the observed signal. These range from techniques interpretable through a linear regression framework (Weaverdyck & Huterer, 2021) such as mode deprojection (also called template deprojection; Rybicki & Press, 1992; Slosar et al., 2004; Elsner et al., 2016, 2017), template subtraction (Ross et al., 2011; Ho et al., 2012b), and the iterative systematics decontamination (ISD) technique employed by the Dark Energy Survey (DES; Elvin-Poole et al., 2018; Rodríguez-Monroy et al., 2022), to more sophisticated methods employing machine learning tools such as artificial neural networks (Rezaie et al., 2020) or self-organising maps (Johnston et al., 2021).